OpenStudy (anonymous):

What is the answer to 2x(x-7) - 2x2 = 14? Is it no solution?

OpenStudy (amoodarya):

distribute 2x by (x-7) then simplify it is easy to say

OpenStudy (amoodarya):

2x^2-14x -2x^2=14 2x^2 cancel -14x=14 x=-1

OpenStudy (ikram002p):

convert it to the general form ax^2+bx+c=0 so it goes like this 0x^2+14x+14=0 2x^2-14x -2x^2=14 find the determinant b^2-4ac 14^2-0*14=14^2>0 so it has a solution :)